OpenStudy (anonymous):

If f(x) = the integral from -2 to x^4 of sqrt(t^2+8)dt then f'(x)=? I will write this as an equation as well below

OpenStudy (anonymous):

find f'(x) of \[F(x)=\int\limits_{-2}^{x^4}\sqrt{t^2+8}dt\]

OpenStudy (anonymous):

Sorry It's x^8

OpenStudy (anonymous):

\[\sqrt{x^8+8}\]

OpenStudy (slaaibak):

i'd go with \[4x^3 \times \sqrt{x^8 + 8}\]

OpenStudy (sirm3d):

sub x in f(t), and apply the chain rule to x^4

OpenStudy (anonymous):

thank you so much for the help it is much appreciated :)

OpenStudy (anonymous):

Sorry, I'm wrong, I forgot to account for the upper limit being x^4 not just x